Determine the Concentration of H+ , OH-, pH, and pOH of the following solutions:

1. 0.0100 M HCl
2. 0.0200 M HOCl [ HOCl(aq) + H2O(l) ⇄ H3O+(aq) + OCl-(aq); Ka = 3.0 x 10 -8]
1
Expert's answer
2020-05-27T13:35:10-0400

Solutiom (1):

When HCl molecules dissolve they dissociate into H+ ions and Cl- ions. HCl is a strong acid because it dissociates almost completely.

HCl(aq) = H+(aq) + Cl-(aq)

Therefore,

C(HCl) = [H+] = 0.0100 M

[H+] = 0.01 M

We can convert between [H+] and pH using the following equations:

pH = - log[H+]

pH = - log(0.0100) = 2

pH = 2

For any aqueous solution at 25C:

pH + pOH = 14

pOH = 14 - pH = 14 - 2 = 12

pOH = 12

We can convert between [OH-] and pOH using the following equations:

pOH = - log[OH-]

[OH-] = 10-pOH

[OH-] = 10-12 = 1×10-12 M

[OH-] = 1×10-12 M


Answer (1): [H+] = 0.01 M; pH = 2; pOH = 12; [OH-] = 1×10-12 M.


Solution (2):

Since hypochlorous acid (HOCl) is a weak acid and remains mostly undissociated, the major species in a 0.0200 M HOCl solution are HOCl and H2O.

Both species can produce H+. Since HOCl is a significantly stronger acid than H2O, we can neglect water as a source of H+.

The dominant equilibrium is:

HOCl(aq) + H2O(l) ⇄ H3O+(aq) + OCl-(aq)

or

HOCl(aq) = H+(aq) + OCl-(aq)

We therefore use the following equilibrium expression:



The concentrations are as follows:



[HOCl] = 0.02 - x

[H+] = [OCl-] = x

Then,



Assuming that 0.02 >> x, then:

(0.02 - x) = 0.02

Ka = x2 / 0.02

x2 = (3.0×10-8 × 0.02) = 6×10-10

x = (Ka)0.5 = (6×10-10)0.5 = 2.45×10-5

x= [H+] = 2.45×10-5 M

[H+] = 2.45×10-5 M

We can convert between [H+] and pH using the following equations:

pH = - log[H+]

pH = - log(2.45×10-5) = 4.61

pH = 4.61

For any aqueous solution at 25C:

pH + pOH = 14

pOH= 14 - pH = 14 - 4.61 = 9.39

pOH = 9.39

We can convert between [OH-] and pOH using the following equations:

pOH = - log[OH-]

[OH-] = 10-pOH

[OH-] =10-9.39 = 4.07×10-10 M

[OH-] = 4.07×10-10 M


Answer (2): [H+] = 2.45×10-5 M; pH = 4.61; pOH = 9.39; [OH-] = 4.07×10-10 M.